1.4. Calculate the following
1.4.2. How much will 12 litre juice cost if 1.5 litre costs R24? (3)
1.4.3. Pam poured R240 petrol at a price of R15/litre. How many litres did she pour
1
Expert's answer
2020-08-17T18:08:20-0400

1.4.2. Let "x=" the cost of 12 litre juice. We have a proportion


"\dfrac{x}{12}=\dfrac{R24}{1.5}"

"x=\dfrac{12}{1.5}\cdot R24"

"x=R192"

12 litre juice costs R192.


1.4.3. Let "x=" the number of litres she poured. Use the ratio


"x=\dfrac{R240}{R15/litre}=16\ litres"

Pam poured 16 litres of petrol.